# Sample Complexity of Diffusion Model Training Without Empirical Risk Minimizer Access

### v2

@@ D.2 Bounding the Statistical Error
$\mathbb{E}\left[\left|\mathcal{L}_{k}(\theta)-\widehat{\mathcal{L}}_{t}(\theta)\right|\right]\leq\mathcal{O}\left(d\cdot\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right),~{}~{}\quad\forall~{}\theta\in\Theta^{\prime\prime}$ (65)
@@
$\|\theta^{a}_{t}-\theta^{b}_{t}\|^{2}\leq\mu\left|\mathcal{L}_{k}(\theta^{a}_{t})-\mathcal{L}_{k}(\theta^{b}_{t})\right|,$ (66)
and applying Lipschitz continuity of the score network with respect to parameters:
$\leq\mathcal{O}\left(d.\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right).$ (69)
Here $L$ is the lipschitz parameter of the neural network. Taking expectation with respect to $x\sim p_{t_{k}}$ on both sides completes the proof. For the final bound we replace $\delta$ with $\frac{\delta}{2}$ for notational consistency.
D.3 Bounding Optimization Error
@@
Recall Lemma 3 (Optimization Error.) Let $n$ be the number of samples used to estimate the score function at each diffusion step. If the step size satisfies $0\leq\eta\leq\frac{1}{\kappa}$, then under Assumptions 1, 2 and 3, for all $k\in[0,K]$, the optimization error due to imperfect minimization of the training loss satisfies with probability at least $1-\delta$:
$\mathcal{E}^{\mathrm{opt}}_{k}\leq\mathcal{O}\left(d.\sqrt{\frac{\log\frac{2}{\delta}}{n}}\right).$ (70)
@@
^$\widehat{\mathcal{L}}^{i}_{k}(\theta)=\left\|s_{\theta}(x_{i},t_{k})-\nabla\log p_{t}(x_{i})\right\|^{2}.$ (71)
@@
From the smoothness of $\mathcal{L}_{k}(\theta)$ throgh Assumption 1, we have
@@
$\mathcal{L}_{k}(\theta_{i+1})\leq\mathcal{L}_{k}(\theta_{i})+\langle\nabla\mathcal{L}_{k}(\theta_{i}),\theta_{i+1}-\theta_{i}\rangle+\frac{\kappa}{2}\|\theta_{i+1}-\theta_{i}\|^{2}.$ (73)
Taking conditional expectation given $\theta_{i}$, and using the unbiased-ness of the stochastic gradient ^$\nabla\widehat{\mathcal{L}}_{k}(\theta_{i})$, we get:
$\mathbb{E}[\mathcal{L}_{k}(\theta_{i+1})\mid\theta_{i}]\leq\mathcal{L}_{k}(\theta_{i})-\alpha_{t}\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}+\frac{\kappa\alpha^{2}_{t}}{2}\mathbb{E}[\|\nabla\widehat{\mathcal{L}}_{t}(\theta_{i})\|^{2}\mid\theta_{i}].$ (74)
Now using the variance bound on the stochastic gradients using Assumption 3, we have
$\mathbb{E}[\|\nabla\widehat{\mathcal{L}}_{k}(\theta_{i})\|^{2}\mid\theta_{i}]\leq\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}+\sigma^{2},$ (75)
Using this in the previous equation, we have that
@@
$\mathbb{E}[\mathcal{L}_{k}(\theta_{t+1})\mid\theta_{i}]$ $\leq\mathcal{L}_{k}(\theta_{i})-\eta\|\nabla\mathcal{L}(\theta_{i})\|^{2}+\frac{\kappa\eta^{2}}{2}\left(\|\nabla\mathcal{L}(\theta_{i})\|^{2}+\sigma^{2}\right)$ (76)
@@
$=\mathcal{L}(\theta_{i})-\left(\eta-\frac{\kappa\eta^{2}}{2}\right)\|\nabla\mathcal{L}(\theta_{i})\|^{2}+\frac{\kappa\eta^{2}\sigma^{2}}{2}.$ (77)
Now applying the PL inequality (Assumption 2), $\|\nabla L(\theta_{i})\|^{2}\geq 2\mu\left(L(\theta_{i})-L^{*}\right)$, we substitute in the above inequality to get

### v5

@@ C.1 Bounding the Statistical Error
$\left|\mathcal{L^{\prime}}_{k}(\theta)-\widehat{\mathcal{L^{\prime}}}_{k}(\theta)\right|\leq\mathcal{O}\left(\frac{d\cdot W^{D}}{n}\right)+\mathcal{O}\left(\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right),~~\quad\forall~\theta\in\Theta^{\prime\prime}.$ (48)
This yields that with probability at least $1-\delta$ we have
$\left|\mathcal{L^{\prime}}_{k}(\theta)-\widehat{\mathcal{L^{\prime}}}_{k}(\theta)\right|\leq\mathcal{O}\left(d\cdot W^{D}\cdot\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right),~~\quad\forall~\theta\in\Theta^{\prime\prime}$ (49)
From this we have
$\left[\left|\mathcal{L^{\prime}}_{k}(\theta)-\widehat{\mathcal{L}}_{k}(\theta)\right|\right]\leq\mathcal{O}\left(d\cdot\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right)$ (50)
Now consider the probability of the event
$A_{i,j}=\left\{\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa\right\}$ (51)
Where $\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|_{j}$ denotes the $k^{th}$ co-ordinate of of the $i^{th}$ sample of the score function given by $\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|$ We have the probability of this event upper bounded as
$P\left(\Bigg|\frac{x_{i}-e^{-t}x_{0}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa\right)$ $=\mathbb{E}_{z}P\left(\Bigg|\frac{x_{i}-e^{-t}x_{0}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa\right)$ (52)
$\leq\exp\left(-{\kappa^{2}(1-e^{-t})}\right)$ (53)
$\leq\exp\left(-{\kappa^{2}}\right)$ (54)
Setting $\kappa={{\log}\left(\frac{dn}{\delta}\right)}$, we have
$P\left(\Bigg|\frac{x_{i}-e^{-t}x_{0}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa\right)\leq\frac{\delta}{dn}$ (55)
If we denote the event $A=\{\widehat{L}^{\prime}(\theta)=\widehat{L}(\theta)\}$, then by union bound we have $P(A)=P(\cup_{i,j}A_{i,j})\leq\sum_{i,j}P(A_{i,j})\leq\delta$. Let event $B$ denote the failure of the generalization bound, i.e.,
$B:=\left\{\left|\mathcal{L^{\prime}}_{t}(\theta)-\widehat{\mathcal{L^{\prime}}}_{t}(\theta)\right|>\widehat{R}(\Theta^{\prime\prime})+\mathcal{O}\left(\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right)\right\}.$ (56)
From above, we know $\mathbb{P}(B)\leq\delta$ under the boundedness condition. Therefore, by the union bound, we have
$\mathbb{P}(A\cup B)$ $\leq\mathbb{P}(A)+\mathbb{P}(B)\leq 2\delta,$ (57)
$\implies\mathbb{P}(A^{c}\cap B^{c})$ $=1-P(A\cup B)\geq 1-2\delta.$ (58)
On this event $(A^{c}\cap B^{c})$, we have $\widehat{\mathcal{L}}^{\prime}(\theta)=\widehat{\mathcal{L}}(\theta)$. Hence, with probability at least $1-2\delta$, we have
$|\mathcal{L}_{k}(\theta_{k}^{b})-\mathcal{L}_{k}(\theta_{k}^{a})|$ $\leq\left|\mathcal{L}_{k}(\theta_{k}^{b})-\widehat{\mathcal{L}}_{t}(\theta_{k}^{b})\right|+\left|\mathcal{L}_{k}(\theta_{k}^{a})-\widehat{\mathcal{L}}_{t}(\theta_{k}^{a})\right|.$ (59)
$\leq\left|\mathcal{L}_{k}(\theta_{k}^{b})-\widehat{\mathcal{L^{\prime}}}_{t}(\theta_{k}^{b})\right|+\left|\mathcal{L}_{k}(\theta_{k}^{a})-\widehat{\mathcal{L^{\prime}}}_{t}(\theta_{k}^{a})\right|.$ (60)
$=\left|\mathcal{L}_{k}(\theta_{k}^{b})-{\mathcal{L^{\prime}}}_{t}(\theta_{k}^{b})\right|+\left|\mathcal{L}_{k}(\theta_{k}^{a})-{\mathcal{L^{\prime}}}_{t}(\theta_{k}^{a})\right|.$
$\quad+\left|\mathcal{L^{\prime}}_{t}(\theta_{k}^{b})-\widehat{\mathcal{L^{\prime}}}_{t}(\theta_{k}^{b})\right|+\left|\mathcal{L}_{k}(\theta_{k}^{a})-\widehat{\mathcal{L^{\prime}}}_{t}(\theta_{k}^{a})\right|.$ (61)
$\leq\left|\mathcal{L^{\prime}}_{t}(\theta_{k}^{a})-{\mathcal{L}}_{t}(\theta_{k}^{a})\right|+\left|\mathcal{L^{\prime}}_{t}(\theta_{k}^{b})-{\mathcal{L}}_{t}(\theta_{k}^{b})\right|+\mathcal{O}\left(d\cdot W^{D}\cdot\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right)$ (62)
In order to bound $|\mathcal{L}_{k}(\theta)-\mathcal{L^{\prime}}_{t}(\theta)|$ we have the following
$|\mathcal{L}_{k}(\theta)-\mathcal{L^{\prime}}_{t}(\theta)|$ $\leq\sum_{j=1}^{d}\mathbb{E}_{x_{j}\sim(u_{t_{k}})_{j}}|(s_{\theta}(x,t_{k}))-({\nabla}{\log}p_{t_{k}}(x))|_{j}^{2}-\mathbb{E}_{x\sim u_{t}}|(v_{t}(x))_{k}-(v_{\theta}(x,t))|_{j}^{2}$ (63)
$\leq\sum_{j=1}^{d}\mathbb{E}_{x_{j}\sim(u_{t_{k}})_{j}}\left(|({\nabla}{\log}p_{t_{k}})-(s_{\theta}(x,t_{k}))|_{j}^{2}\textbf{1}_{\Bigg|\frac{x-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)$ (64)
$\leq\sum_{j=1}^{d}\mathbb{E}_{x_{j}\sim(u_{t_{k}})}\left(\Bigg|\frac{x-e^{-t}(x_{0})}{\sigma_{t}^{2}}-(s_{\theta}(x,t))_{k}\Bigg|_{j}^{2}\textbf{1}_{\Bigg|\frac{x-e^{-t}(x_{0})}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)$ (65)
$\leq 2\sum_{j=1}^{d}\mathbb{E}_{x_{j}\sim(u_{t_{k}})_{j}}\left(\Bigg|\frac{x-e^{-t}(x_{0})}{\sigma_{t}^{2}}\Bigg|_{j}^{2}\textbf{1}_{\Bigg|\frac{x-e^{-t}(x_{0})}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)$
$+\sum_{j=1}^{d}2\mathbb{E}_{x_{j}\sim(u_{t_{k}})_{j}}\left((s_{\theta}(x,t))_{j}^{2}\textbf{1}_{\Bigg|\frac{x-e^{-t}(x_{0})}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)$ (66)
$\leq 2\sum_{j=1}^{d}\mathbb{E}_{x_{j}\sim(u_{t_{k}})_{j}}\left(\Bigg|\frac{x-e^{-t}(x_{0})}{\sigma_{t}^{2}}\Bigg|_{j}^{2}\textbf{1}_{\Bigg|\frac{x-e^{-t}(x_{0})}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)$
$+\sum_{j=1}^{d}C_{\Phi^{{}^{\prime\prime}}}\mathbb{E}_{x_{j}\sim(u_{t_{k}})_{j}}\left(\Bigg|\frac{x-e^{-t}(x_{0})}{\sigma_{t}^{2}}\Bigg|_{j}\textbf{1}_{\Big|\frac{x-e^{-t}(x_{0})}{\sigma_{t}^{2}}\Big|\geq\kappa}\right)$ (67)
$\leq\left(\frac{4+2C_{\Phi^{{}^{\prime\prime}}}}{{\sigma_{t}}^{2}}\right)\underbrace{\sum_{k=1}^{d}\mathbb{E}_{x_{j}\sim(u_{t_{k}})_{j}}\left(|x|_{j}^{2}\textbf{1}_{\Bigg|\frac{x-e^{-t}(x_{0})}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)}_{I}$
$+\frac{2}{{\sigma_{t}}^{2}}\underbrace{\mathbb{E}_{x_{j}\sim(u_{t_{k}})_{j}}\left(|x_{0}|_{j}^{2}\textbf{1}_{\Bigg|\frac{x-e^{-t}(x_{0})}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)}_{II}$ (68)
We get Equation equation 66 from Equation equation 65 by using the identity $(a-b)^{2}\leq 2|a|^{2}+2|b|^{2}$. We get Equation equation 67 from Equation equation 66 by using Lemma 8. We get Equation equation 68 from Equation equation 67 by using the identity $(a-b)^{2}\leq 2|a|^{2}+2|b|^{2}$ again. Now we separately obtain upper bounds for the terms $I$ and $II$ as follows.
$\sum_{j=1}^{d}\mathbb{E}_{x_{j}\sim(u_{t_{k}})_{j}}\left(|x_{j}|^{2}\textbf{1}_{\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)$ (69)
$=\sum_{j=1}^{d}\mathbb{E}_{x_{j}\sim(u_{t_{k}})_{j}}\left(|x_{j}|^{2}\textbf{1}_{\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)$ (70)
$\leq\sum_{j=1}^{d}\mathbb{E}_{x_{0}}\mathbb{E}_{x_{j}\sim(u_{t_{k}})_{j}|x_{0}}\left(|x_{j}|^{2}\textbf{1}_{\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)$ (71)
$\leq\sum_{j=1}^{d}\mathbb{E}_{x_{0}}\mathbb{E}_{x_{j}\sim(u_{t})_{k}|x_{0}}\left(|x_{j}|^{2}\Bigg|\textbf{1}_{\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)P\left(\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|\geq\kappa\Bigg|x_{0}\right)$ (72)
$\leq{\exp}\left(-{\kappa}^{2}\right)\sum_{j=1}^{d}\mathbb{E}_{x_{0}}\mathbb{E}_{x_{0}\sim(u_{t})_{j}|x_{0}}\left(|x_{k}|^{2}\Bigg|_{j}\textbf{1}_{\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|\geq\kappa}\right)$ (73)
$\leq{\exp}\left(-{\kappa}^{2}\right)\sum_{k=1}^{d}\mathbb{E}_{x_{0}}\left(\sigma^{2}_{t}+\sigma^{2}_{t}.{\kappa{\cdot}\sigma^{2}_{t}}.\frac{{\phi}(\kappa{\cdot}\sigma^{2}_{t})}{1-\Phi((\kappa{\cdot}\sigma^{2}_{t}))}\right)$ (74)
$\leq{\exp}\left(-{\kappa}^{2}\right)\sum_{k=1}^{d}\mathbb{E}_{z}\left(2.\sigma^{2}_{t}\right)$ (75)
$\leq\mathcal{O}\left({\exp}\left(-{\kappa}^{2}\right)\right)$ (76)
We get Equation equation 74 from Equation equation 73 by using Lemma 7. We get Equation equation 76 from Equation equation 75 from Assumption 1 and by using the upper bound on the Mill’s ration which implies that $\frac{{\phi}(\kappa)}{1-\Phi(\kappa)}\leq{\kappa}+\frac{1}{\kappa}$. We get Equation equation 76 from Equation equation 75 from Assumption 1, which implies that the second moment of $z$ is bounded.
$\sum_{j=1}^{d}\mathbb{E}_{x_{j}\sim(u_{t_{k}})_{j}}\left(|x|_{0}^{2}\textbf{1}_{\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)$ (77)
$=\sum_{j=1}^{d}\mathbb{E}_{x_{k}\sim(u_{t})_{k}}\left(|x_{0}|^{2}\textbf{1}_{\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)$ (78)
$\leq\sum_{j=1}^{d}\mathbb{E}_{x_{0}}\mathbb{E}_{x_{k}\sim(u_{t})_{k}|x_{0}}\left(|x_{0}|^{2}\textbf{1}_{\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)$ (79)
$\leq\sum_{j=1}^{d}\mathbb{E}_{x_{0}}\mathbb{E}_{x_{k}\sim(u_{t})_{k}|x_{0}}\left(|x_{0}|^{2}\Bigg|\textbf{1}_{\Bigg|\frac{x_{i}-e^{-t}(x_{0})_{i}}{\sigma_{t}^{2}}\Bigg|_{j}\geq\kappa}\right)P\left(\Big|\frac{x_{k}-tz_{k}}{1-t}\Big|_{j}\geq\kappa|x_{0}\right)$ (80)
$\leq{\exp}\left(-{\kappa}^{2}\right)\sum_{j=1}^{d}\mathbb{E}_{x_{0}}|x_{0}|^{2}$ (81)
$\leq\mathcal{O}\left({\exp}\left(-{\kappa}^{2}\right)\right)$ (82)
Setting $\kappa=\log\frac{dn}{\delta}$ Plugging Equation equation 76, equation 82 into Equation equation 68. Then we have
$|\mathcal{L}_{k}(\theta)-\mathcal{L^{\prime}}_{t}(\theta)|\leq$ $\mathcal{O}\left({\exp}\left(-{\kappa}^{2}\right)\right),\quad\forall\theta=\{\theta_{k}^{a},\theta_{k}^{b}\}$ (83)
$\leq$ $\mathcal{O}\left(\frac{\delta}{dn}\right)$ (84)
Now plugging Equation equation 83 into Equation equation 62 we get with probability at least $1-2{\delta}$
$|\mathcal{L}_{k}(\theta_{k}^{b})-\mathcal{L}_{k}(\theta_{k}^{a})|$ $\leq\mathcal{O}\left(d\cdot W^{D}\cdot\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right)$ (85)
@@
$\|\theta_{k}^{a}-\theta_{k}^{b}\|^{2}\leq\mu\left|\mathcal{L}_{k}(\theta_{k}^{a})-\mathcal{L}_{k}(\theta_{k}^{b})\right|,$ (86)
and applying Lipschitz continuity of the velocity fields with respect to parameter $x$
$\|v^{\theta^{a}_{k}}(x,t_{k})-v^{\theta^{b}_{k}}(x,t_{k})\|^{2}$ $\leq L_{t}\cdot\|\theta_{k}^{a}-\theta_{k}^{b}\|^{2}$ (87)
$\leq L_{t}\cdot\mu\left|\mathcal{L}_{k}(\theta_{k}^{a})-\mathcal{L}_{k}(\theta_{k}^{b})\right|$ (88)
$\leq\mathcal{O}\left(d\cdot W^{D}\cdot\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right).$ (89)
Here $L_{t}$ is the Lipschitz parameter of the neural networks. It is always possible to obtain this Lipschitz constant as the quantity $\|v^{a}(x,t)-v^{b}(x,t)\|^{2}\leq L_{t}$ is non-zero only over a finite domain of $x$. Taking expectation with respect to $x$, we obtain the following.
$\mathbb{E}_{x\sim u_{t_{k}}}\mathbb{\|}s^{\theta^{a}_{k}}(x,t_{k})-s^{\theta^{b}_{k}}(x,t_{k})\|^{2}$ $\leq 2\mathbb{E}_{x\sim u_{t}}\mathbb{\|}s^{\theta^{a}_{k}}(x,t_{k})-s^{\theta^{b}_{k}}(x,t_{k})-v^{a}_{t}(x)-v^{b}_{t}(x)\|^{2}$
$+2\mathbb{E}_{x\sim u_{t}}\mathbb{\|}v^{\theta^{a}_{k}}(x,t_{k})-v^{\theta^{b}_{k}}(x,t_{k})\|^{2}$ (90)
$\leq 4\mathbb{E}_{x\sim u_{t}}\mathbb{\|}v^{\theta^{a}_{k}}(x,t_{k})-s^{\theta^{a}_{k}}(x,t_{k})\mathbb{\|}^{2}$ (91)
$+4{\mathbb{E}_{x\sim u_{t}}}\mathbb{\|}s^{\theta^{b}_{k}}(x,t_{k})-v^{\theta^{b}_{k}}(x,t_{k})\|^{2}$
$+4\mathbb{E}_{x\sim u_{t}}\mathbb{\|}v^{a}_{t}(x)-v^{b}_{t}(x)\|^{2}$ (92)
$\leq\mathcal{O}(\kappa^{-2})+\mathcal{O}\left(d\cdot W^{D}\cdot\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right).$ (93)
$\leq\mathcal{O}(\frac{\delta}{dn})+\mathcal{O}\left(d\cdot W^{D}\cdot\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right).$ (94)
$\leq\mathcal{O}\left(d\cdot W^{D}\cdot\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right).$ (95)
This completes the proof. Note that the quantities $4\mathbb{E}_{x\sim u_{t}}\mathbb{\|}u^{a}_{t}(x)-v^{a}_{t}(x)\mathbb{\|}^{2}$ and $4\mathbb{E}_{x\sim u_{t}}\mathbb{\|}u^{a}_{t}(x)-v^{a}_{t}(x)\mathbb{\|}^{2}$ are bounded in the same manner as is done in Equation equation 82.
C.2 Bounding Optimization Error
@@
$\widehat{\mathcal{L}}^{i}_{k}(\theta)=\left\|s_{\theta}(x_{i},t_{k})-\nabla\log p_{t}(x_{i})\right\|^{2}.$ (96)
@@
Consider stochastic gradient descent (SGD) iterates indexed by $i=1,2,\dots,n$:
$\theta_{i+1}\;=\;\theta_{i}-\eta_{i}\,\nabla\hat{\mathcal{L}}(\theta_{i}),$
with diminishing stepsizes
$\eta_{i}=\frac{\alpha}{i+\gamma},\qquad\alpha>0,\ \gamma>0,$ (98)
chosen so that
$\alpha\mu>1\quad\text{and}\quad\eta_{i}\leq\frac{1}{L}\ \ \text{for all }i\;\;(\text{e.g.\ take }\gamma\geq\alpha L).$ (99)
Define the expected suboptimality $e_{i}:=\mathbb{E}\bigl[\mathcal{L}_{k}(\theta_{i})-\mathcal{L}_{k}^{\star}\bigr]$.
By Assumption 4 for $\mathcal{L}_{k}$ gives, for $y=\theta_{i+1}=\theta_{i}-\eta_{i}\nabla\hat{\mathcal{L}_{k}}(\theta_{i})$,
$\mathcal{L}_{k}(\theta_{i+1})$ $\leq\mathcal{L}_{k}(\theta_{i})+\Big\langle\nabla\mathcal{L}_{k}(\theta_{i}),\theta_{i+1}-\theta_{i}\Big\rangle+\frac{L}{2}\|\theta_{i+1}-\theta_{i}\|^{2}$
@@
$=\mathcal{L}_{k}(\theta_{i})-\eta_{i}\Big\langle\nabla\mathcal{L}_{k}(\theta_{i}),\nabla\hat{\mathcal{L}_{k}}(\theta_{i})\Big\rangle+\frac{L}{2}\eta_{i}^{2}\bigl\|\nabla\hat{\mathcal{L}_{k}}(\theta_{i})\bigr\|^{2}.$
Taking conditional expectation given $\theta_{i}$ and using unbiasedness and the variance bound,
$\mathbb{E}\!\bigl[\|\nabla\hat{\mathcal{L}_{k}}(\theta_{i})\|^{2}\mid\theta_{i}\bigr]=\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}+\mathbb{E}\!\bigl[\|\nabla\hat{\mathcal{L}_{k}}(\theta_{i})-\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}\mid\theta_{i}\bigr]\leq\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}+\sigma^{2},$
we obtain
@@
$\mathbb{E}\!\left[\mathcal{L}_{k}(\theta_{i+1})\mid\theta_{i}\right]\leq\mathcal{L}_{k}(\theta_{i})-\eta_{i}\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}+\frac{L}{2}\eta_{i}^{2}\Bigl(\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}+\sigma^{2}\Bigr).$ (100)
If $\eta_{i}\leq\frac{1}{L}$ then $\eta_{i}+\tfrac{L}{2}\eta_{i}^{2}\leq-\tfrac{\eta_{i}}{2}$, hence
@@
$\mathbb{E}\!\left[\mathcal{L}_{k}(\theta_{i+1})\mid\theta_{i}\right]\leq\mathcal{L}_{k}(\theta_{i})-\frac{\eta_{i}}{2}\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}+\frac{L}{2}\eta_{i}^{2}\sigma^{2}.$ (101)
By Assumption 2, we have that $\|\nabla\mathcal{L}_{k}(\theta_{i})\|^{2}\geq 2\mu\,(\mathcal{L}_{k}(\theta_{i})-\mathcal{L}_{k}^{\star})$, so